I would like help finding the equation to this problem please:

The sum of two numbers is 20. Find the numbers if their product is 96.

I have gotten the first part (I think). It is x+y=20. I just do not know how to add the second part of the equation to the first part.
Thank you for your help!! :)

Answer by scott8148(6628)

good, so far

x = 20 - y

x * y = 96

substituting ___ (20 - y) y = 96 ___ 20y - y^2 = 96 ___ 0 = y^2 - 20y + 96

factoring ___ 0 = (y - 12)(y - 8)

y = 12 ___ or ___ y = 8 ___ and x = 8 ___ or ___ x = 12

the two numbers are 8 and 12